A bag contains 4 yellow,3 pink,2 red, and 1 blue marbles. A marble is drawn at random with replacement. Find each probability.
P(yellow,pink)
P(blue, blue)
P(red,pink,blue)
Thankyou

Answer by Fombitz(32388)

4 Y (yellow), 3 P (pink), 2 R (red), 1 B (blue)=10 total
.
.
.
P(Y)=4/10=2/5
P(P)=3/10
P(Y,P)=(2/5)(3/10)=3/25
.
.
.
P(B)=1/10
P(B,B)=(1/10)(1/10)=1/100
.
.
.
P(R)=2/10=1/5
P(P)=3/10
P(B)=1/10
P(R,P,B)=(1/5)(3/10)(1/10)=3/500